OpenStudy (macabbey101):

HELP WILL FAN AND MEDAL!:) Write the equation of a line that is perpendicular to the given line and that passes through the given point. y + 5 = −5/7 (x – 1); (1, –5)

OpenStudy (igreen):

Perpendicular lines have opposite slope, flip the fraction and multiply by -1 to find the slope of the new line.

OpenStudy (macabbey101):

Can you explain a little more please?

OpenStudy (macabbey101):

@iGreen

OpenStudy (igreen):

I think your equation is missing something..check back at the question..is there an 'x' in it?

OpenStudy (macabbey101):

it says 4+5=-5/7(x-1) (1,-5)

OpenStudy (igreen):

Oh wait, nevermind. y + 5 = -5/7(x - 1) This equation is in point-slope form: y - y1 = m(x - x1) Where 'm' is the slope and (x1, y1) is a point on the line.

OpenStudy (macabbey101):

okay. thanks

OpenStudy (igreen):

? What did you come up with?

OpenStudy (igreen):

The slope here is -5/7, so flip the fraction and multiply by 1 to find the perpendicular line's slope.

OpenStudy (igreen):

multiply by -1*

OpenStudy (macabbey101):

so you would do -7/5*-1?

OpenStudy (igreen):

Correct!

OpenStudy (macabbey101):

would that be for all perpendicular equations or just this one in particular?:)

OpenStudy (macabbey101):

@iGreen

OpenStudy (macabbey101):

@iGreen i got 7/5... so how would I write that for my answer? like how would i put that in equation form

OpenStudy (igreen):

Okay, we have the point and the slope, so all we have to do is plug it back into point-slope form. \(\sf y - y_1 = m(x - x_1)\) Where \(\sf m\) is the slope and \(\sf (x_1, y_1)\) is a point on the line. Just plug in the given information..we already found the slope, and the point is given in the question.

OpenStudy (macabbey101):

so i would write it as y-7/5x-1? I think thats right..?

OpenStudy (macabbey101):

y+7/5x-1 sorry messed up there

OpenStudy (macabbey101):

ugh thats supposed to be an equals sign sorry lol

OpenStudy (igreen):

Does it want the equation in slope-intercept form or point-slope form? Just plug in the values..don't change anything else.

OpenStudy (igreen):

\(\sf y - y_1 = m(x - x_1)\) Where \(\sf m\) is the slope and \(\sf (x_1, y_1)\) is a point on the line. In this case: \(\sf m = \dfrac{7}{5}\) And: \(\sf x_1 = 1\) \(\sf y_1 = -5\)

OpenStudy (igreen):

So what do we have?